A sample of 330 with x-bar = 75 and Sx = 11 has a standard error equal to what?

A. .033

B. .227

C. Cannot be determined since z* is not given

D. 5.477

E. .606

F. 4.129

Find the critical t* value for a two-sided test at the α = 0.01 level based on the t distribution where n = 13.

A. 2.650

B. 2.576

C. 2.681

D. 2.326

E. 3.012

F. 3.055
Answer #1

Given:

n = 330

= 75

S = 11

Therefore,

Standard error (SE) =

ANSWER: E

E. 0.606

Question 12.

n = 13

= 0.01

Two sided test

Degree of freedom = n-1 = 13-1 = 12

Critical value:

        ................Using standard Normal table

ANSWER: F

F. 3.055